\begin{thebibliography}{44}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aepli et~al.(2023)Aepli, {\c{C}}{\"o}ltekin, van~der Goot,
  Jauhiainen, Kazzaz, Ljube{\v{s}}i{\'c}, North, Plank, Scherrer, and
  Zampieri}]{2023-findings-vardial}
No{\"e}mi Aepli, {\c{C}}a{\u{g}}r{\i} {\c{C}}{\"o}ltekin, Rob van~der Goot,
  Tommi Jauhiainen, Mourhaf Kazzaz, Nikola Ljube{\v{s}}i{\'c}, Kai North,
  Barbara Plank, Yves Scherrer, and Marcos Zampieri. 2023.
\newblock Findings of the {V}ar{D}ial evaluation campaign 2023.
\newblock In \emph{Proceedings of the Tenth Workshop on NLP for Similar
  Languages, Varieties and Dialects}, Dubrovnik, Croatia. Association for
  Computational Linguistics.

\bibitem[{Bandel et~al.(2022)Bandel, Aharonov, Shmueli-Scheuer, Shnayderman,
  Slonim, and Ein-Dor}]{bandel-etal-2022-quality}
Elron Bandel, Ranit Aharonov, Michal Shmueli-Scheuer, Ilya Shnayderman, Noam
  Slonim, and Liat Ein-Dor. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.acl-long.45} {Quality
  controlled paraphrase generation}.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 596--609,
  Dublin, Ireland. Association for Computational Linguistics.

\bibitem[{Bastianelli et~al.(2020)Bastianelli, Vanzo, Swietojanski, and
  Rieser}]{bastianelli-etal-2020-slurp}
Emanuele Bastianelli, Andrea Vanzo, Pawel Swietojanski, and Verena Rieser.
  2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.588} {{SLURP}: A
  spoken language understanding resource package}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 7252--7262, Online. Association
  for Computational Linguistics.

\bibitem[{Bilat et~al.(2020)Bilat, Dinu, and Cieliebak}]{bilat2020cross}
Lo{\"i}s Bilat, Georgiana Dinu, and Mark Cieliebak. 2020.
\newblock Cross-lingual toxicity detection.
\newblock pages 50--55.

\bibitem[{{Caswell} et~al.(2021){Caswell}, {Kreutzer}, {Wang}, {Wahab}, {van
  Esch}, {Ulzii-Orshikh}, {Tapo}, {Subramani}, {Sokolov}, {Sikasote},
  {Setyawan}, {Sarin}, {Samb}, {Sagot}, {Rivera}, {Rios}, {Papadimitriou},
  {Osei}, {Ortiz Su{\'a}rez}, {Orife}, {Ogueji}, {Niyongabo}, {Nguyen},
  {M{\"u}ller}, {M{\"u}ller}, {Hassan Muhammad}, {Muhammad}, {Mnyakeni},
  {Mirzakhalov}, {Matangira}, {Leong}, {Lawson}, {Kudugunta}, {Jernite},
  {Jenny}, {Firat}, {Dossou}, {Dlamini}, {de Silva}, {{\c{C}}abuk Ball{\i}},
  {Biderman}, {Battisti}, {Baruwa}, {Bapna}, {Baljekar}, {Abebe Azime},
  {Awokoya}, {Ataman}, {Ahia}, {Ahia}, {Agrawal}, and
  {Adeyemi}}]{caswell-etal-2021-quality}
Isaac {Caswell}, Julia {Kreutzer}, Lisa {Wang}, Ahsan {Wahab}, Daan {van Esch},
  Nasanbayar {Ulzii-Orshikh}, Allahsera {Tapo}, Nishant {Subramani}, Artem
  {Sokolov}, Claytone {Sikasote}, Monang {Setyawan}, Supheakmungkol {Sarin},
  Sokhar {Samb}, Beno{\^\i}t {Sagot}, Clara {Rivera}, Annette {Rios}, Isabel
  {Papadimitriou}, Salomey {Osei}, Pedro~Javier {Ortiz Su{\'a}rez}, Iroro
  {Orife}, Kelechi {Ogueji}, Rubungo~Andre {Niyongabo}, Toan~Q. {Nguyen},
  Mathias {M{\"u}ller}, Andr{\'e} {M{\"u}ller}, Shamsuddeen {Hassan Muhammad},
  Nanda {Muhammad}, Ayanda {Mnyakeni}, Jamshidbek {Mirzakhalov}, Tapiwanashe
  {Matangira}, Colin {Leong}, Nze {Lawson}, Sneha {Kudugunta}, Yacine
  {Jernite}, Mathias {Jenny}, Orhan {Firat}, Bonaventure F.~P. {Dossou},
  Sakhile {Dlamini}, Nisansa {de Silva}, Sakine {{\c{C}}abuk Ball{\i}}, Stella
  {Biderman}, Alessia {Battisti}, Ahmed {Baruwa}, Ankur {Bapna}, Pallavi
  {Baljekar}, Israel {Abebe Azime}, Ayodele {Awokoya}, Duygu {Ataman},
  Orevaoghene {Ahia}, Oghenefego {Ahia}, Sweta {Agrawal}, and Mofetoluwa
  {Adeyemi}. 2021.
\newblock \href {http://arxiv.org/abs/2103.12028} {{Quality at a Glance: An
  Audit of Web-Crawled Multilingual Datasets}}.
\newblock \emph{arXiv e-prints}, page arXiv:2103.12028.

\bibitem[{Chen et~al.(2019)Chen, Zhuo, and
  Wang}]{DBLP:journals/corr/abs-1902-10909}
Qian Chen, Zhu Zhuo, and Wen Wang. 2019.
\newblock \href {http://arxiv.org/abs/1902.10909} {{BERT} for joint intent
  classification and slot filling}.
\newblock \emph{CoRR}, abs/1902.10909.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and
  Stoyanov}]{conneau-etal-2020-unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised
  cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 8440--8451, Online. Association for
  Computational Linguistics.

\bibitem[{Conneau and Lample(2019)}]{lample2019cross}
Alexis Conneau and Guillaume Lample. 2019.
\newblock \href
  {https://proceedings.neurips.cc/paper/2019/file/c04c19c2c2474dbf5f7ac4372c5b9af1-Paper.pdf}
  {Cross-lingual language model pretraining}.
\newblock \emph{Advances in neural information processing systems}, 32.

\bibitem[{Coucke et~al.(2018)Coucke, Saade, Ball, Bluche, Caulier, Leroy,
  Doumouro, Gisselbrecht, Caltagirone, Lavril, Primet, and
  Dureau}]{coucke2018snips}
Alice Coucke, Alaa Saade, Adrien Ball, Th{\'e}odore Bluche, Alexandre Caulier,
  David Leroy, Cl{\'e}ment Doumouro, Thibault Gisselbrecht, Francesco
  Caltagirone, Thibaut Lavril, Ma{\"e}l Primet, and Joseph Dureau. 2018.
\newblock \href {http://arxiv.org/abs/1805.10190} {Snips voice platform: an
  embedded spoken language understanding system for private-by-design voice
  interfaces}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin2019bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://www.aclweb.org/anthology/N19-1423.pdf} {{BERT:
  Pre-training of Deep Bidirectional Transformers for Language Understanding}}.
\newblock In \emph{Proceedings of the 2019 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186.

\bibitem[{Feng et~al.(2020{\natexlab{a}})Feng, Yang, Cer, Arivazhagan, and
  Wang}]{feng2020languageagnostic}
Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang.
  2020{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/2007.01852} {Language-agnostic bert
  sentence embedding}.

\bibitem[{Feng et~al.(2020{\natexlab{b}})Feng, Yang, Cer, Arivazhagan, and
  Wang}]{feng2020language}
Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang.
  2020{\natexlab{b}}.
\newblock Language-agnostic bert sentence embedding.
\newblock \emph{arXiv preprint arXiv:2007.01852}.

\bibitem[{FitzGerald et~al.(2022)FitzGerald, Hench, Peris, Mackie, Rottmann,
  Sanchez, Nash, Urbach, Kakarala, Singh, Ranganath, Crist, Britan, Leeuwis,
  Tur, and Natarajan}]{https://doi.org/10.48550/arxiv.2204.08582}
Jack FitzGerald, Christopher Hench, Charith Peris, Scott Mackie, Kay Rottmann,
  Ana Sanchez, Aaron Nash, Liam Urbach, Vishesh Kakarala, Richa Singh, Swetha
  Ranganath, Laurie Crist, Misha Britan, Wouter Leeuwis, Gokhan Tur, and Prem
  Natarajan. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2204.08582} {Massive: A
  1m-example multilingual natural language understanding dataset with 51
  typologically-diverse languages}.

\bibitem[{Frey et~al.(2016)Frey, Glaznieks, and Stemle}]{inproceedings}
Jennifer Frey, Aivars Glaznieks, and Egon Stemle. 2016.
\newblock The didi corpus of south tyrolean cmc data: A multilingual corpus of
  facebook texts.

\bibitem[{Goo et~al.(2018)Goo, Gao, Hsu, Huo, Chen, Hsu, and
  Chen}]{goo2018slot}
Chih-Wen Goo, Guang Gao, Yun-Kai Hsu, Chih-Li Huo, Tsung-Chieh Chen, Keng-Wei
  Hsu, and Yun-Nung Chen. 2018.
\newblock Slot-gated modeling for joint slot filling and intent prediction.
\newblock In \emph{Proceedings of The 16th Annual Conference of the North
  American Chapter of the Association for Computational Linguistics: Human
  Language Technologies}.

\bibitem[{Gupta et~al.(2018)Gupta, Shah, Mohit, Kumar, and
  Lewis}]{gupta-etal-2018-semantic-parsing}
Sonal Gupta, Rushin Shah, Mrinal Mohit, Anuj Kumar, and Mike Lewis. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1300} {Semantic parsing for
  task oriented dialog using hierarchical representations}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 2787--2792, Brussels, Belgium.
  Association for Computational Linguistics.

\bibitem[{Hu et~al.(2021)Hu, Shen, Wallis, Allen{-}Zhu, Li, Wang, and
  Chen}]{DBLP:journals/corr/abs-2106-09685}
Edward~J. Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen{-}Zhu, Yuanzhi Li,
  Shean Wang, and Weizhu Chen. 2021.
\newblock \href {http://arxiv.org/abs/2106.09685} {Lora: Low-rank adaptation of
  large language models}.
\newblock \emph{CoRR}, abs/2106.09685.

\bibitem[{Li et~al.(2021)Li, Arora, Chen, Gupta, Gupta, and
  Mehdad}]{li-etal-2021-mtop}
Haoran Li, Abhinav Arora, Shuohui Chen, Anchit Gupta, Sonal Gupta, and Yashar
  Mehdad. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.eacl-main.257} {{MTOP}: A
  comprehensive multilingual task-oriented semantic parsing benchmark}.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  2950--2962, Online. Association for Computational Linguistics.

\bibitem[{Li et~al.(2018)Li, Wang, Sun, Panda, Liu, and Gao}]{li2018microsoft}
Xiujun Li, Yu~Wang, Siqi Sun, Sarah Panda, Jingjing Liu, and Jianfeng Gao.
  2018.
\newblock \href {http://arxiv.org/abs/1807.11125} {Microsoft dialogue
  challenge: Building end-to-end task-completion dialogue systems}.

\bibitem[{Linder et~al.(2020)Linder, Jungo, Hennebert, Musat, and
  Fischer}]{linder2020crawler}
Lucy Linder, Michael Jungo, Jean Hennebert, Claudiu~Cristian Musat, and Andreas
  Fischer. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.329} {Automatic
  creation of text corpora for low-resource languages from the internet: The
  case of swiss german}.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 2706--2711, Marseille, France. European Language Resources
  Association.

\bibitem[{Liu and Lane(2016)}]{liu2016attention}
Bing Liu and Ian~R. Lane. 2016.
\newblock \href {http://arxiv.org/abs/1609.01454} {Attention-based recurrent
  neural network models for joint intent detection and slot filling}.
\newblock \emph{CoRR}, abs/1609.01454.

\bibitem[{Liu et~al.(2019)Liu, Eshghi, Swietojanski, and
  Rieser}]{liu2019benchmarking}
Xingkun Liu, Arash Eshghi, Pawel Swietojanski, and Verena Rieser. 2019.
\newblock \href {http://arxiv.org/abs/1903.05566} {Benchmarking natural
  language understanding services for building conversational agents}.

\bibitem[{Lugosch et~al.(2019)Lugosch, Ravanelli, Ignoto, Tomar, and
  Bengio}]{lugosch2019speech}
Loren Lugosch, Mirco Ravanelli, Patrick Ignoto, Vikrant~Singh Tomar, and Yoshua
  Bengio. 2019.
\newblock \href {https://doi.org/10.21437/Interspeech.2019-2396} {{Speech Model
  Pre-Training for End-to-End Spoken Language Understanding}}.
\newblock In \emph{Proc. Interspeech 2019}, pages 814--818.

\bibitem[{Mangrulkar et~al.(2022)Mangrulkar, Gugger, Debut, Belkada, and
  Paul}]{peft}
Sourab Mangrulkar, Sylvain Gugger, Lysandre Debut, Younes Belkada, and Sayak
  Paul. 2022.
\newblock Peft: State-of-the-art parameter-efficient fine-tuning methods.
\newblock \url{https://github.com/huggingface/peft}.

\bibitem[{Mesnil et~al.(2015)Mesnil, Dauphin, Yao, Bengio, Deng, Hakkani-Tur,
  He, Heck, Tur, Yu, and Zweig}]{6998838}
Grégoire Mesnil, Yann Dauphin, Kaisheng Yao, Yoshua Bengio, Li~Deng, Dilek
  Hakkani-Tur, Xiaodong He, Larry Heck, Gokhan Tur, Dong Yu, and Geoffrey
  Zweig. 2015.
\newblock \href {https://doi.org/10.1109/TASLP.2014.2383614} {Using recurrent
  neural networks for slot filling in spoken language understanding}.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}, 23(3):530--539.

\bibitem[{Muennighoff et~al.(2022)Muennighoff, Wang, Sutawika, Roberts,
  Biderman, Scao, Bari, Shen, Yong, Schoelkopf
  et~al.}]{muennighoff2022crosslingual}
Niklas Muennighoff, Thomas Wang, Lintang Sutawika, Adam Roberts, Stella
  Biderman, Teven~Le Scao, M~Saiful Bari, Sheng Shen, Zheng-Xin Yong, Hailey
  Schoelkopf, et~al. 2022.
\newblock Crosslingual generalization through multitask finetuning.
\newblock \emph{arXiv preprint arXiv:2211.01786}.

\bibitem[{{NLLB Team} et~al.(2022){NLLB Team}, Costa-jussà, Cross, Çelebi,
  Elbayad, Heafield, Heffernan, Kalbassi, Lam, Licht, Maillard, Sun, Wang,
  Wenzek, Youngblood, Akula, Barrault, Mejia-Gonzalez, Hansanti, Hoffman,
  Jarrett, Sadagopan, Rowe, Spruit, Tran, Andrews, Ayan, Bhosale, Edunov, Fan,
  Gao, Goswami, Guzmán, Koehn, Mourachko, Ropers, Saleem, Schwenk, and
  Wang}]{nllb2022}
{NLLB Team}, Marta~R. Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad,
  Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht,
  Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al~Youngblood, Bapi
  Akula, Loic Barrault, Gabriel Mejia-Gonzalez, Prangthip Hansanti, John
  Hoffman, Semarley Jarrett, Kaushik~Ram Sadagopan, Dirk Rowe, Shannon Spruit,
  Chau Tran, Pierre Andrews, Necip~Fazil Ayan, Shruti Bhosale, Sergey Edunov,
  Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn,
  Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and
  Jeff Wang. 2022.
\newblock No language left behind: Scaling human-centered machine translation.

\bibitem[{Price(1990)}]{price-1990-evaluation}
P.~J. Price. 1990.
\newblock \href {https://aclanthology.org/H90-1020} {Evaluation of spoken
  language systems: the {ATIS} domain}.
\newblock In \emph{Speech and Natural Language: Proceedings of a Workshop Held
  at Hidden Valley, {P}ennsylvania, June 24-27,1990}.

\bibitem[{Ramshaw and Marcus(1995)}]{ramshaw-marcus-1995-text}
Lance Ramshaw and Mitch Marcus. 1995.
\newblock \href {https://aclanthology.org/W95-0107} {Text chunking using
  transformation-based learning}.
\newblock In \emph{Third Workshop on Very Large Corpora}.

\bibitem[{Rasley et~al.(2020)Rasley, Rajbhandari, Ruwase, and
  He}]{10.1145/3394486.3406703}
Jeff Rasley, Samyam Rajbhandari, Olatunji Ruwase, and Yuxiong He. 2020.
\newblock \href {https://doi.org/10.1145/3394486.3406703} {Deepspeed: System
  optimizations enable training deep learning models with over 100 billion
  parameters}.
\newblock In \emph{Proceedings of the 26th ACM SIGKDD International Conference
  on Knowledge Discovery \& Data Mining}, KDD '20, page 3505–3506, New York,
  NY, USA. Association for Computing Machinery.

\bibitem[{Ravuri and Stolcke(2015)}]{InterspeechRecurrent}
Suman Ravuri and Andreas Stolcke. 2015.
\newblock \href {https://doi.org/10.21437/Interspeech.2015-42} {Recurrent
  neural network and lstm models for lexical utterance classification}.
\newblock pages 135--139.

\bibitem[{Reimers and Gurevych(2019)}]{reimers2019sentence}
Nils Reimers and Iryna Gurevych. 2019.
\newblock Sentence-bert: Sentence embeddings using siamese bert-networks.
\newblock \emph{arXiv preprint arXiv:1908.10084}.

\bibitem[{Reimers and Gurevych(2020)}]{reimers2020making}
Nils Reimers and Iryna Gurevych. 2020.
\newblock Making monolingual sentence embeddings multilingual using knowledge
  distillation.
\newblock \emph{arXiv preprint arXiv:2004.09813}.

\bibitem[{Schuster et~al.(2019)Schuster, Gupta, Shah, and
  Lewis}]{schuster-etal-2019-cross-lingual}
Sebastian Schuster, Sonal Gupta, Rushin Shah, and Mike Lewis. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1380} {Cross-lingual transfer
  learning for multilingual task oriented dialog}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 3795--3805,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Tiedemann(2012)}]{tiedemann2012parallel}
J{\"o}rg Tiedemann. 2012.
\newblock Parallel data, tools and interfaces in opus.
\newblock In \emph{Lrec}, volume 2012, pages 2214--2218. Citeseer.

\bibitem[{Tiedemann and Thottingal(2020)}]{TiedemannThottingal:EAMT2020}
J{\"o}rg Tiedemann and Santhosh Thottingal. 2020.
\newblock {OPUS-MT} — {B}uilding open translation services for the {W}orld.
\newblock In \emph{Proceedings of the 22nd Annual Conferenec of the European
  Association for Machine Translation (EAMT)}, Lisbon, Portugal.

\bibitem[{Upadhyay et~al.(2018)Upadhyay, Faruqui, T{\"u}r, Dilek, and
  Heck}]{8461905}
Shyam Upadhyay, Manaal Faruqui, Gokhan T{\"u}r, Hakkani-T{\"u}r Dilek, and
  Larry Heck. 2018.
\newblock \href {https://doi.org/10.1109/ICASSP.2018.8461905} {(almost)
  zero-shot cross-lingual spoken language understanding}.
\newblock In \emph{2018 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 6034--6038.

\bibitem[{van~der Goot et~al.(2021)van~der Goot, Sharaf, Imankulova,
  {\"U}st{\"u}n, Stepanovi{\'c}, Ramponi, Khairunnisa, Komachi, and
  Plank}]{van-der-goot-etal-2021-masked}
Rob van~der Goot, Ibrahim Sharaf, Aizhan Imankulova, Ahmet {\"U}st{\"u}n,
  Marija Stepanovi{\'c}, Alan Ramponi, Siti~Oryza Khairunnisa, Mamoru Komachi,
  and Barbara Plank. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.197} {From masked
  language modeling to translation: Non-{E}nglish auxiliary tasks improve
  zero-shot spoken language understanding}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 2479--2497, Online. Association for Computational
  Linguistics.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/file/3f5ee243547dee91fbd053c1c4a845aa-Paper.pdf}
  {Attention is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  6000--6010.

\bibitem[{Xu and Sarikaya(2013)}]{xu2013convolutional}
Puyang Xu and Ruhi Sarikaya. 2013.
\newblock \href {https://doi.org/10.1109/ASRU.2013.6707709} {Convolutional
  neural network based triangular crf for joint intent detection and slot
  filling}.
\newblock pages 78--83.

\bibitem[{Xu et~al.(2020)Xu, Haider, and Mansour}]{xu-etal-2020-end}
Weijia Xu, Batool Haider, and Saab Mansour. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.410} {End-to-end
  slot alignment and recognition for cross-lingual {NLU}}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 5052--5063, Online. Association
  for Computational Linguistics.

\bibitem[{Xue et~al.(2020)Xue, Constant, Roberts, Kale, Al-Rfou, Siddhant,
  Barua, and Raffel}]{xue2020mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya
  Siddhant, Aditya Barua, and Colin Raffel. 2020.
\newblock mt5: A massively multilingual pre-trained text-to-text transformer.
\newblock \emph{arXiv preprint arXiv:2010.11934}.

\bibitem[{Yao et~al.(2014)Yao, Peng, Zweig, Yu, Li, and Gao}]{6854368}
Kaisheng Yao, Baolin Peng, Geoffrey Zweig, Dong Yu, Xiaolong Li, and Feng Gao.
  2014.
\newblock \href {https://doi.org/10.1109/ICASSP.2014.6854368} {Recurrent
  conditional random field for language understanding}.
\newblock In \emph{2014 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 4077--4081.

\bibitem[{Zhang et~al.(2021)Zhang, Hashimoto, Wan, Liu, Xiong, and
  Yu}]{DBLP:journals/corr/abs-2106-04564}
Jianguo Zhang, Kazuma Hashimoto, Yao Wan, Ye~Liu, Caiming Xiong, and Philip~S.
  Yu. 2021.
\newblock \href {http://arxiv.org/abs/2106.04564} {Are pretrained transformers
  robust in intent classification? {A} missing ingredient in evaluation of
  out-of-scope intent detection}.
\newblock \emph{CoRR}, abs/2106.04564.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EACL2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage{footnote}
\usepackage{array}
\usepackage{graphicx}
\usepackage{color, colortbl}
\usepackage{xstring}
\usepackage{comment}
\usepackage{adjustbox}
\usepackage{float}
\restylefloat{table}
\usepackage{array,booktabs,makecell}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{listings}
\usepackage{xcolor}

\usepackage{appendix}
\usepackage{times,latexsym}
\usepackage{url}
\usepackage[T1]{fontenc}
\usepackage{lipsum}  
\usepackage{graphicx}
\usepackage{xcolor} 
\usepackage{soul}
\usepackage{listings}
\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\colorlet{punct}{red!60!black}
\definecolor{background}{HTML}{EEEEEE}
\definecolor{delim}{RGB}{20,105,176}
\colorlet{numb}{magenta!60!black}

\lstdefinelanguage{json}{
    basicstyle=\normalfont\ttfamily,
    stepnumber=1,
    showstringspaces=false,
    breaklines=true,
    frame=lines,
    backgroundcolor=\color{white},
    literate=
     *{0}{{{\color{numb}0}}}{1}
      {1}{{{\color{numb}1}}}{1}
      {2}{{{\color{numb}2}}}{1}
      {3}{{{\color{numb}3}}}{1}
      {4}{{{\color{numb}4}}}{1}
      {5}{{{\color{numb}5}}}{1}
      {6}{{{\color{numb}6}}}{1}
      {7}{{{\color{numb}7}}}{1}
      {8}{{{\color{numb}8}}}{1}
      {9}{{{\color{numb}9}}}{1}
      {:}{{{\color{punct}{:}}}}{1}
      {,}{{{\color{punct}{,}}}}{1}
      {\{}{{{\color{delim}{\{}}}}{1}
      {\}}{{{\color{delim}{\}}}}}{1}
      {[}{{{\color{delim}{[}}}}{1}
      {]}{{{\color{delim}{]}}}}{1},
}

\title{Zero-Shot Slot and Intent Detection in Low-Resource Languages}

\author{\normalsize Sang Yun Kwon$^{1,\star}$ ~ Gagan Bhatia $^{1,\star}$ ~ El Moatez Billah Nagoudi$^{1}$ 
~\\ {\bf Alcides Alcoba Inciarte}$^{1}$ ~{\bf Muhammad Abdul-Mageed}$^{1,2}$\\
\normalsize $^{1}$Deep Learning \& Natural Language Processing Group,
  The University of British Columbia\\\normalsize  $^{2}$Department of Natural Language Processing \& Department of Machine Learning, MBZUAI\\ %
  \texttt{\normalsize \{skwon01@student.,gagan30@student.,moatez.nagoudi@,alcobaaj@mail.,muhammad.mageed@\}ubc.ca}}

\begin{document}
\maketitle
\begin{abstract}
Intent detection and slot filling are critical tasks in spoken and natural language understanding for task-oriented dialog systems. In this work we describe our participation in the slot and intent detection for low-resource language varieties  (SID4LR;~\citet{2023-findings-vardial}).  We investigate the slot and intent  detection 
 (SID) tasks using a wide range of models and settings. Given the recent success of multitask-prompted finetuning of large language models, we also test the generalization capability of the recent encoder-decoder model mT0~\cite{muennighoff2022crosslingual} on new tasks (i.e., SID)  in languages they have never intentionally seen. We show that our best model outperforms the baseline by a large margin (up to $+30$ F\textsubscript{1} points) in both SID tasks.

\end{abstract}
\section{Introduction}\label{sec:intro}


Digital conversational assistants have become increasingly pervasive. Examples of popular virtual assistants include Siri, Alexa, and Google. A crucial factor in the effectiveness of these systems is their capacity to understand user input and respond or act accordingly to fulfill particular requirements. Most of these applications are voice-based and hence need spoken language understanding (SLU). SLU typically starts with automatic speech recognition (ASR), taking the sound of spoken language and transcribing it into text. Then, it handles natural language understanding (NLU) tasks to extract semantic features from the text including question answering, dialogue management,  intent detection, and slot filling.

The intent detection task aims to recognize the speaker's desired outcome from a given utterance. And slot filling focuses on identifying the main arguments or the spans of words in the utterance that contain semantic information relevant to the intent. Table~\ref{tab:xsid_ex} shows four utterances in different languages: English, Swiss German (GSW), South Tyrolean (ST), and Neapolitan (NAP). The English example has the intent \textit{set\_alarm} and two individual spans \textit{Set an alarm} and \textit{6 am on Wed} are labeled with their slot tags \textit{location} and \textit{datetime}, respectively, using the Inside, Outside, Beginning (IOB) \cite{ramshaw-marcus-1995-text} tagging format.

\begin{table}
\centering
 \renewcommand{\arraystretch}{1.75}
\resizebox{1\columnwidth}{!}{%

\begin{tabular}{ll}
 \toprule
\textbf{Lang.} & \textbf{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Annotation} \\  \toprule
\textbf{EN} & \colorbox{green!35}{Set an alarm} for\colorbox{orange!35}{6 am on Wed} \\
\textbf{GSW } & Du em \colorbox{green!35}{Mittwuch e Wecker} dry fürem sächsi em \colorbox{orange!35}{Morge}. \\ 
\textbf{ST} & Stell an \colorbox{green!35}{Wecker} firn \colorbox{orange!35}{Mittig af 6 in der friah} \\
\textbf{NAP} & Imposta 'na \colorbox{green!35}{sveglia} 'e \colorbox{orange!35}{6 'e matina 'e miercurì} \\ \toprule
\end{tabular}}
\caption{Examples of xSID annotations in our target languages from the validation set with intents (alarm / set\_alarm) and slots (\colorbox{green!35}{location}, \colorbox{orange!35}{datetime}). \textbf{EN}: English, \textbf{GSW}:Swiss German \textbf{ST}: South Tyrolean, \textbf{NAP}: Neapolitan} \label{tab:xsid_ex} 
\end{table}

In this work, we present our participation in the  slot and intent detection for low-resource language varieties (SID4LR;~\citet{2023-findings-vardial}) shared task. The shared task takes as its target  
three low resources languages-- Swiss German (GSW), South Tyrolean (ST), and Neapolitan (NAP).  
The main objective of the SID4LR shared task is to find the most effective approach for transferring knowledge to less commonly spoken languages that have limited resources and lack a standard writing system, in the zero-shot setting (i.e., without use of any training data). In the context of the shared task, we target the following four main research questions:

\begin{enumerate}

    \item[\textbf{Q1:}] Can successful models on English SID tasks be generalizable to new unseen languages (i.e., the zero-shot setting)?

      \item[\textbf{Q2:}] How do models trained  on a language from the given language family fare on a low-resource variety from the same family under the zero-shot setting (i.e., with no access to training data from these low-resource varieties). For example, in our case, we ask how do models trained on German perform on Swiss German or South Tyrolean, and how do models trained on Italian perform on Neapolitan. 

\item[\textbf{Q3:}] What impact does exploiting data augmentation techniques such as  paraphrasing and machine translation have on the SID tasks in the zero-shot context?

   \item[\textbf{Q4:}] Are the existing large multilingual models, trained using multitask-prompted fine-tuning, able to achieve zero-shot generalization to SID tasks in languages that they have never intentionally seen?


\end{enumerate}

The rest of this paper is organized as follows: Section~\ref{sec:RW} is a literature review on intent and slot  detection tasks. The shared task, the source data provided in SID4LR, and  the external parallel data we exploit to build our models  are described in Section ~\ref{sec:SID4LR}. In Section~\ref{sec:data}, we provide information about datasets, baselines, and data preprocessing. The baseline, and multilingual pre-trained language models we used are described in Section~\ref{sec:models}. We present our experimental settings and our training procedures in Section~\ref{sec:exp}. Section~\ref{sec:results} is an analysis and discussion of our 
results.  And we conclude in Section~\ref{sec:conc}.

\section{Related Work}\label{sec:RW}

The problem of low-resource slot and intent detection for languages with limited training data has been the focus of several recent research works. In this section, we discuss some of the most relevant and recent works, including datasets, benchmarks, and models that aim to address this challenge.

\subsection{\textbf{SID Benchmarks and Corpus}}

The table below provides an overview of various datasets used for NLU tasks. These datasets cover a range of languages, domains, intents, and slots, and are widely used to evaluate the performance of NLU models. Some of the prominent datasets include MASSIVE, SLURP, NLU Evaluation Data, ATIS, MultiATIS++, Snips, TOP, MTOP, Cross-lingual Multilingual Task-Oriented Dialog, Microsoft Dialog Challenge, and Fluent Speech Commands. These datasets have been used for tasks such as intent classification, slot filling, and semantic parsing. Overall, these datasets provide a useful resource for researchers to benchmark their models and develop better NLU systems.

\begin{table*}[!ht]
\resizebox{\textwidth}{!}{%
\begin{tabular}{lccccc}
\toprule
\textbf{Name}                                         & \# \textbf{Langs} & \textbf{Utt. per Lang (K)} & \textbf{Domains} & \textbf{Intents} & \textbf{Slots}                                                                                                           \\ \midrule
Airline Travel Information System (ATIS) \citep{price-1990-evaluation} &$1$ &$5.8$ &$1$ &$26$ &$129$ \\
ATIS with Hindi and Turkish \citep{8461905} &$3$ &$1.3$-$5.8$ &$1$ &$26$ &$129$ \\
Cross-lingual Multilingual Task Oriented Dialog \citep{schuster-etal-2019-cross-lingual} &$3$ &$5.08$-$43.3$ &$3$ &$12$ &$11$ \\
Fluent Speech Commands (FSC) \citep{lugosch2019speech} &$1$ &$30$ &- &$31$ &- \\
MASSIVE \cite{https://doi.org/10.48550/arxiv.2204.08582} &$51$ &$19.5$ &$18$ &$60$ &$55$ \\
Microsoft Dialog Challenge \citep{li2018microsoft} &1 &$38.2$ &3 &11 &29 \\
MultiATIS++ \citep{xu-etal-2020-end} &$9$ &$1.4$-$5.8$ &$1$ &$21$-$26$ &$99$-$140$ \\
Multilingual Task-Oriented Semantic Parsing (MTOP) \citep{li-etal-2021-mtop} &$6$ &$15.1$-$22.2$ &$11$ &$104$-$113$ &$72$-$75$ \\
NLU Evaluation Data \citep{liu2019benchmarking} &$1$ &$25,7$ &$18$ &$54$ &$56$ \\
SLURP \citep{bastianelli-etal-2020-slurp} &1 &$16,5$ &18 &60 &55 \\
SNIPS \citep{coucke2018snips} &$1$ &$14.4$ &- &$7$ &$53$ \\
Task Oriented Parsing (TOP) \citep{gupta-etal-2018-semantic-parsing} &$1$ &$44.8$ &$2$ &$25$ &$36$ \\
xSID \cite{van-der-goot-etal-2021-masked} &$13$ &$10$ &$7$ &$16$ &$33$ \\

\bottomrule
\end{tabular}}
\caption{  SID benchmark and datasets with the number of languages covered, number of utterances per language, domain, intent count, and slot count. }\label{table:NLUDatasets}
\end{table*}

\subsection{SID Approaches and Models}

The are many works devoted to the SID tasks. Most of these works are categorized into three approaches: \textit{(1) single model for intent detection, (2)  single model for  slot filling, and (3) joint model}.

\noindent\textbf{(1) Single Model for Intent Detection} refers to developing a single model that can identify the intent behind a user's spoken or written input. This approach involves training a neural network or other machine learning model on a large dataset of labeled examples. Each example consists of user input and its corresponding intent label. The model then uses this training data to learn patterns and features that can accurately predict the intent of new user inputs. For instance,  \newcite{InterspeechRecurrent} proposed 
a recurrent neural network and LSTM models for intent detection in spoken language understanding. In this work, the authors first discuss the limitations of traditional intent detection approaches that rely on handcrafted features and propose using deep learning models to learn features directly from the data.  \newcite{DBLP:journals/corr/abs-2106-04564} investigate the robustness of pre-trained transformers-based models such as BERT and RoBERTa for intent classification in spoken language understanding. They conduct experiments on two datasets, ATIS~\cite{8461905} and SNIPS~\cite{coucke2018snips}, showing that pre-trained transformers perform well on in-scope intent detection.

\noindent \textbf{(2) Single Model for Slot Filling} is an approach that aims to develop a single model capable of identifying slots in spoken language understanding. The model takes a sentence as input and predicts the slot labels for each word in the sentence. Various recurrent neural network (RNN) architectures such as Elman-type \cite{6998838} and Jordan-type \cite{6998838} networks and their variants have been explored to find the most effective architecture for slot filling. Incorporating word embeddings has also been studied and found to improve slot-filling performance significantly.  For example, \newcite{6854368} use LSTM networks with word embeddings for slot filling on the ATIS~\cite{8461905} dataset and achieve state-of-the-art (SOTA) results at the time.
\newcite{goo2018slot} propose a bi-directional LSTM (BLSTM) with an attention mechanism for slot filling on the  ATIS~\cite{8461905} and SNIPS~\cite{coucke2018snips} datasets.

\noindent\textbf{(3) Joint Model} is an approach that aims to jointly model the intent detection and slot-filling tasks in spoken language understanding. This approach trains a single model to predict both the intent and slot labels simultaneously. The model uses the context of the input sentence to predict these labels. Joint models have been shown to achieve SOTA performance on several spoken language understanding datasets. \newcite{xu2013convolutional} propose a joint convolutional neural network (CNN) and RNN model for intent detection and slot filling on the ATIS~\cite{8461905} dataset. They achieved SOTA results at the time. In the same context, \newcite{liu2016attention} proposed an attention-based neural network for joint intent detection and slot filling. The model uses an attention mechanism to weigh the importance of different parts of the input sentence for predicting the intent label and slot labels. \newcite{DBLP:journals/corr/abs-1902-10909} explore the use of the BERT model for joint intent detection and slot filling on ATIS~\cite{8461905} and SNIPS~\cite{coucke2018snips}. They report SOTA results on both datasets.

\section{SID4LR Shared Task}\label{sec:SID4LR}

\noindent{\textbf{Task Formulation.}} Intent detection and slot-filling are critical NLP tasks where, given an utterance, a system is responsible for parsing the user's intent and extracting relevant information to act or reply appropriately. While many neural-based models have achieved SOTA performance for these tasks, their success often depends on large amounts of labeled data. However, many real-world datasets are limited to specific domains and are only available in English or a few other languages. As a result, it is important to reuse existing data in high-resource languages to develop models for low-resource languages, especially since tasks like intent classification and slot-filling require abundant labeled data. 

\noindent{\textbf{Shared Task Problem Statement.} This shared task of SID aims to address the challenges of performing SID for low-resource language varieties for the following languages: Swiss German, South Tyrolean, and Neapolitan. 
The training data provided consists of the Cross-lingual Slot and Intent Detection (xSID\textsubscript{0.4}) corpus \cite{van-der-goot-etal-2021-masked}, a cross-lingual spoken language understanding dataset, covering $12$ languages (Arabic, Chinese, Dutch, Danish, English, German, Indonesian, Italian, Japanese, Kazakh, Serbian, Turkish) from six language families with English training. The task allowed the use of pre-trained models and external data including data from the target language.}

\noindent{\bf{Evaluation Metric.}} The primary evaluation metric for slot filling is the span F\textsubscript{1} score, where both span and label must match exactly, and accuracy is used to evaluate intent detection where it is calculated through the ratio of the number of correct predictions of intent to the total number of sentences. More details regarding the shared task can be found in  \citet{2023-findings-vardial}.

\section{Data}\label{sec:data}

\noindent\textbf{Shared Task Data.} The xSID\textsubscript{0.4} \cite{van-der-goot-etal-2021-masked} corpus comprises cross-lingual SLU evaluation datasets covering $13$ languages from six language families. The training dataset contains $43,605$ sentences, the development set contains $300$ sentences, and the test set contains $500$ sentences. The corpus contains sentences from Snips and Facebook, which were translated into all $13$ target languages, resulting in a cross-lingual SLU evaluation dataset. All examples are annotated with their intent and corresponding slots. Tables~\ref{tab:data1} and ~\ref{tab:dat2} provides examples of annotations with intent and slots. We converted the dataset into a JSON format that includes intents and slots. A sample of the resulting JSON format is show below.

\begin{table}
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{lrrr}
\toprule
\textbf{Language} &  \# \textbf{Train} &  \# \textbf{Valid} &  \# \textbf{Test} \\
\midrule
   ar &  $42,157$ &    $300$ &   $500$ \\
   da &  $43,605$ &    $300$ &   $500$ \\
   de &  $43,605$ &    $300$ &   $500$ \\
   en &  $43,605$ &    $300$ &   $500$ \\

   id &  $42,157$ &    $300$ &   $500$ \\
   it &  $43,605$ &    $300$ &   $500$ \\
   ja &  $29,073$ &    $150$ &   $250$ \\
   kk &  $42,157$ &    $300$ &   $500$ \\
   nl &  $43,605$ &    $300$ &   $500$ \\
   sr &  $43,605$ &    $300$ &   $500$ \\
   tr &  $43,605$ &    $300$ &   $500$ \\
   zh &  $42,157$ &    $300$ &   $500$ \\
\bottomrule
\end{tabular}}
\caption{ Number of samples in the train, validation, and test sets for each language in the  multilingual dataset xSID\textsubscript{0.4}, where the language codes are represented by two-letter ISO codes. The dataset includes 12 languages: Arabic (ar), Danish (da), German (de), English (en), Indonesian (id), Italian (it), Japanese (ja), Kazakh (kk), Dutch (nl), Serbian (sr), Turkish (tr), and Chinese (zh).}\label{tab:bAbI}
\end{table}

\begin{table}[!htp]\centering
\caption{Example of data summary}\label{tab: }
\scriptsize
\resizebox{1\columnwidth}{!}{%
\begin{tabular}{lrr}\toprule
\# \textcolor{blue}{\texttt{text}}: & \textbf{show all reminders} \\[0.2cm]
\# \textcolor{blue}{\texttt{intent}}: &\texttt{reminder/show\_reminders} \\[0.2cm]
\# \textcolor{blue}{\texttt{slots}}: & \texttt{5:8:reminder/reference, 9:18:reminder/noun} \\
\bottomrule
\end{tabular}}
\end{table}\label{tab:data1}

\begin{table}[!htp]\centering
\caption{In-depth data example}\label{tab: }
\scriptsize
\resizebox{1\columnwidth}{!}{%
\begin{tabular}{lrrr}\toprule
1 \textcolor{red}{show} &reminder/show\_reminders &O \\[0.1cm]
2 \textcolor{red}{all} &reminder/show\_reminders &B-reference \\[0.1cm]
3 \textcolor{red}{reminders} &reminder/show\_reminders &O \\
\bottomrule
\end{tabular}}
\end{table}\label{tab:dat2}

\begin{lstlisting}[language=json,firstnumber=1,caption=Example of preprocessed data,captionpos=b]
{
    'text':'show all reminders',
    'slots': 'reference:all',
    'intent': 'reminder/show_reminders',
    '__index_level_0__': 0
 }
\end{lstlisting}
\label{data:json}

\noindent\textbf{External Data.} As mentioned, Swiss German, South Tyrolean, and Neapolitan are low-resource languages with limited available labeled data. To address this challenge, we incorporate unlabeled data from different sources to augment our training data. We describe these external sources next.

\noindent\textbf{SwissCrawl} \cite{linder2020crawler}, a corpus of over $500,000$ Swiss German sentences gathered from web crawling between September and November $2019$. The sentences are representative of how native speakers write in forums and social media and may contain slang and ascii emojis.

\noindent \textbf{DiDi Corpus} \cite{inproceedings} is a multilingual language corpus of $600,000$ tokens from Facebook users in South Tyrol, Italy. It includes CMC texts, socio-demographic data, and linguistic annotations on thread, text, and token level. The corpus is mainly German and Italian, with English also present, and has been manually anonymized and annotated.

\noindent\textbf{OSCAR Corpus} \cite{caswell-etal-2021-quality} is a large multilingual corpus created by scraping the web and includes texts in more than $200$ languages. The OSCAR Corpus includes texts in Neapolitan, which is a Romance language spoken in the southern part of Italy, particularly in the region of Campania. The Neapolitan texts in the corpus consist of around $4.4$ million tokens, making it one of the largest resources available for this language.

\section{Pre-trained Language Models}\label{sec:models}
In this study, we evaluate several popular multilingual Transformer-based language models, including mBERT, XLM-R, SBERT, LaBSE, LASER, and mT0. These models are capable of effectively capturing cross-lingual embeddings, enabling transfer learning across multiple languages. Below we provide a description of each model used in our experiments on the training dataset. 

   \noindent\textit{\textbf{mBERT.}} is the multilingual version of BERT~\cite{devlin2019bert}, which is an encoder model with bidirectional representations from Transformers trained with a denoising objective.  mBERT is trained on Wikipedia for $104$ languages including German and Italian.

 \noindent\textit{\textbf{XLM-R.}}~\cite{conneau-etal-2020-unsupervised} is a transformer-based multilingual masked language model pre-trained on more than $2$TB  of filtered CommonCrawl data in $100$ languages, including languages including German and Italian. XLM-R uses a Transformer model
~\cite{vaswani2017attention} trained with a multilingual masked language model XLM \cite{lample2019cross}.

\noindent{\textbf{sBERT.}} Sentence-BERT (SBERT)~\cite{reimers2019sentence}, is a modification of the pretrained BERT~\cite{devlin2019bert}  model that uses siamese and triplet network structures to derive semantically meaningful sentence embeddings that can be compared using cosine-similarity. As we work under a multilingual context, we use the  multilingual versions from previously monolingual SBERT models~\cite{reimers2020making} which is trained for sentence embedding in $50$+ languages from various language families.

\noindent{\textbf{LaBSE.}} Language-agnostic BERT Sentence Encoder (LaBSE)~\cite{feng2020languageagnostic} is a BERT-based model  trained to generate sentence embeddings in $109$ different languages.  The model's pre-training approach involves a combination of masked language modeling and translation language modeling. The pre-training process combines masked language modeling with translation language modeling. LaBSE is useful for producing sentence embeddings in multiple languages and performing bi-text retrieval.

\noindent{\textbf{LASER.}} Language-Agnostic Sentence Representations (LASER)~\cite{feng2020language} is a contextualized language model based on a BiLSTM encoder trained on parallel data from OPUS website~\cite{tiedemann2012parallel} using a translation objective. The LASER model can handle $200$ different languages.

\noindent{\textbf{mT0.}} \cite{muennighoff2022crosslingual} is a group of sequence-to-sequence models that come with different sizes from $300$M to $13$B  parameters trained to investigate the cross-lingual generalization through multitask finetuning. mT0  can execute human instructions in many languages without any prior training. The models are fine-tuned from pre-existing mT5 ~\cite{xue2020mt5} multilingual language models using a cross-lingual task mixture called xP3. These refined models are capable of cross-lingual generalization to unseen languages.

\section{Experiments and Settings}\label{sec:exp}

\noindent{\textbf{Training on English Data.}} As a baseline setting, we train all the pre-trained models described in Section~\ref{sec:models} on the English part of the multilingual dataset  xSID\textsubscript{$0.4$}~\cite{van-der-goot-etal-2021-masked}  and evaluate them on Swiss German, South Tyrolean, and Neapolitan under a zero-shot setting.


\noindent{\textbf{Training on German/Italian Data.}} Our second approach aims to train all the pre-trained models on the language family of low-resource languages (i.e., German for Swiss German and South Tyrolean, and Italian for Neapolitan, respectively) under the zero-shot setting. So, we extract the German and Italian SID data from xSID\textsubscript{$0.4$}, and then fine-tune all our models on both datasets. Then, we evaluate the German models on GSW and ST tasks and the Italian models on the NAP task.



\noindent{\textbf{Training on Multilingual Dataset.}} Next, we explore a third training approach that involves the full multilingual xSID\textsubscript{$0.4$} dataset. To do so, we combine all the $12$ available languages in the xSID\textsubscript{$0.4$} dataset and fine-tune our pre-trained models on this combined dataset. We then evaluate each target using a zero-shot setting. This approach allows us to train on larger and more diverse datasets. In total, we generate $502,936$ training sentences across all languages in the dataset. 

\noindent{\textbf{Paraphrase and Machine Translation.}} To improve the performance of our pretrained models, we also explore the impact of data augmentation techniques such as paraphrasing and machine translation. Specifically, we aime to examine how these techniques can enhance the performance of our models on cross-lingual SLU tasks. To this end, we experiment with different data augmentation strategies, including paraphrasing and machine translation. Paraphrasing is performed using the quality-guided controlled paraphrase generation (QCPG) model~\cite{bandel-etal-2022-quality}, resulting in a total of $130,815$ sentences in English. These sentences are then translated into German and Italian using the OPUS-MT model \cite{TiedemannThottingal:EAMT2020}, creating cross-lingual datasets for our experiment. 

To further augment our training data for low-resource languages, we leverage Meta AI's No Language Left Behind (NLLB), which provides open-source models capable of high-quality translations between $200$ languages (including low-resource languages \cite{nllb2022}). To create our new training data using the NLLB model, we first use FastText to detect the language codes of our target languages. Next, we utilize NLLB models to translate the English training data into the predicted language codes. The language codes identified for our target languages are \textit{deu{\_}latn} for Swiss German, \textit{est{\_}latn} for South Tyrolean, and \textit{ita{\_}Latn} for Neapolitan. We generate $43,605$ sentences for each of the three languages. It is worth noting that we ensure that the labels for each sentence remain the same throughout the paraphrasing and machine translation process to maintain the integrity of the data.

\noindent{\textbf{Training on External Data.}
Since the language models we employ do not have a strong representation of the low-resource languages used on the task, we leverage large corpora of each of the low-resource languages into the training process. By incorporating external datasets, the models are exposed to more comprehensive information about the semantics of each low-resource language, enabling them to better capture the nuances and complexities of the target languages.

\noindent{\textbf{Training MT0}}
As discussed in Section \ref{sec:models}, the MT0 models share the same architecture as MT5/T5 models, i.e., they are encoder-decoder models. Therefore, we train them for intent classification and slot detection using the data preprocessing approach described in Section \ref{sec:data}. We utilize the PEFT library provided by Huggingface \cite{peft} to train the mT0-small, mT0-Base, and mt0-Large models. Our approach involves using LORA \cite{DBLP:journals/corr/abs-2106-09685}, which allows us to achieve SOTA performance while consuming significantly less memory. For the mT0-xxl models, we utilize DeepSpeed \cite{10.1145/3394486.3406703} with CPU offloading to train a model with 13B parameters on a 40GB A100 GPU. 

\noindent{\textbf{Combining Models.}}} 
In recent studies, joint learning techniques that combine multiple classification approaches have produced promising results \cite{bilat2020cross}. These approaches involve concatenating the outputs of individual models and passing the resulting output through multiple neural network layers, allowing the resulting network to be trained jointly. In this part of our experiments, we investigate the effectiveness of this approach in zero-shot settings by combining multilingual models. Specifically, we combine LASER embeddings, from the LASER model, with other multilingual models including mBERT, sBERT, LaBSE and XLMR.
\section{Results and Discussions}\label{sec:results}
\begin{table*}[t]\centering
\resizebox{\textwidth}{!}{%
\begin{tabular}{llccccccccc}\toprule
\textbf{Setting} &\textbf{Lang.} &\textbf{mBert} &\textbf{LS} &\textbf{LL} &\textbf{LX} &\textbf{mT0-small} &\textbf{mT0-base} &\textbf{mT0-large} &\textbf{mT0-xxl} \\\midrule

\multirow{3}{*}{\textbf{English}} &GSW &$51.67$ &$45.30$ &$52.70$ &$48.30$ &$69.20$ &$70.20$ &$69.00$ &\underline{80.00} \\
&ST &$61.00$ &$58.00$ &$66.70$ &$61.70$ &$74.50$ &$76.20$ &$79.10$ &\underline{89.00} \\
&NAP &$61.00$ &$55.30$ &$56.00$ &$67.0$ &$71.30$ &$72.00$ &$75.00$ &\underline{76.33} \\\hline

\multirow{2}{*}{\textbf{German}} &GSW &$59.00$ &$74.00$ &$68.00$ &$80.70$ &$69.30$ &$73.33$ &$80.33$ &\underline{84.33} \\
&ST &$59.70$ &$55.70$ &$59.00$ &$51.00$ &$83.33$ &$88.33$ &$84.66$ & \underline{92.00} \\\hline
\textbf{Italian} &NAP &$65.30$ &$63.30$ &$63.70$ &$55.70$ &$77.66$ &$84.66$ &$83.33$ & \underline{86.00} \\\hline

\multirow{3}{*}{\textbf{Multilingual}} &GSW &$59.70$ &$62.70$ &$59.70$ &$53.30$ &$75.00$ &$76.33$ &$84.00$ &\underline{$89.00$} \\
&ST &$60.70$ &$54.70$ &$58.00$ &$56.30$ &$88.33$ &$85.66$ &$90.66$ &\underline{$94.00$} \\
&NAP &$61.30$ &$55.70$ &$59.00$ &$60.30$ &$82.66$ &$84.66$ &$86.00$ &\underline{$87.00$}  \\\hline

\multirow{3}{*}{\textbf{Paraphrase+MT}} &GSW &$45.30$ &$37.30$ &$58.00$ &$64.00$ &$79.00$ &$83.00$ &$84.33$ &\textbf{91.00} \\
&ST &$61.70$ &$61.30$ &$60.70$ &$60.70$ &$90.66$ &$93.00$ &$90.00$ &\textbf{95.66} \\
&NAP &$63.70$ &$60.00$ &$58.70$ &$60.00$ &$85.66$ &\textbf{89.00} &$87.33$ &$88.33$ \\
\bottomrule
\end{tabular}}
\caption{Accuracy results for intent classification on the validation set.  \textbf{Baseline:} mBERT~\cite{devlin2019bert}. \textbf{LS:} LASER~\cite{feng2020language}+sBERT~\cite{reimers2019sentence}.  \textbf{LL:} LASER+LaBSE~\cite{feng2020languageagnostic}. \textbf{LX:} LASER+XLM-R~\cite{lample2019cross}. \underline{\textbf{Underline}:} Best-performing models for each setting. \textbf{Bold:}~Best F\textsubscript{1} score across all the experiments and settings. }  \label{tab:dev_res}
\end{table*}

\noindent{\textbf{Evaluation on Validation Data.}} We present the accuracy scores of all our models across various settings. Table~\ref{tab:dev_res} presents the evaluation results for 
the intent classification task on the validation set. Our transformer-based models, with different experimental settings, outperform the baseline on all the target languages. For instance, \textbf{mT0-base} outperforms the baseline (mBERT) with an average of $+16.49$, $+22.93$, $+17.90$ for GSW, ST, and NAP, respectively. Notably, our best combination was the mT0-xxl model under the multilingual setting. It achieves the best results of $89.00$, $94.00$, and $87.00$, improving the baseline with $+29.30$, $+33.30$, and $+25.70$ Accuracy point in the three target languages. 

The  results  of the slot filling task on the validation set are shown in Table~\ref{tab: slots}. Our transformer-based models perform better than the baseline across all target languages when tested under different experimental settings. Our best-performing model, mT0-large, achieves the most outstanding results using the Multilingual settings with F\textsubscript{1} scores of $60.30$, $55.00$, and $52.30$ in the three target languages. These results represent a notable improvement over the baseline, with an increase of $+30.88$, $+4.65$, and $+0.90$ F\textsubscript{1} points in the three target languages. 

Our results on the validation data suggest that larger models generally achieve better performance, implying that higher parameter counts result in better cross-lingual and zero-shot setting performance. Moreover, as the mT0 models are fine-tuned from pre-existing mT5 multilingual language models, they are capable of performing cross-lingual generalization on unseen languages. This capability may be a possible reason for the mT0 models outperforming other models in zero-shot settings.

\begin{table}[t]\centering

\centering
 \renewcommand{\arraystretch}{1.1}
\resizebox{1\columnwidth}{!}{%
\begin{tabular}{llrrrrr}\toprule
\textbf{Setting} &\textbf{Lang.} &\textbf{Baseline} &\textbf{mt0-small} &\textbf{mt0-base} &\textbf{mt0-large} \\\midrule
\multirow{3}{*}{English} &GSW &26.23  &25.42  &34.00  &\textbf{40.32 } \\
&ST &44.61  &32.40  &44.00  &\textbf{54.30 } \\
&NAP &48.01  &42.20  &47.90  &\textbf{49.00 } \\\hline
\multirow{3}{*}{Multi-langl} &GSW &29.42  &28.90  &42.30  &\textbf{60.30 } \\
&ST &50.35  &43.40  &53.40  &\textbf{55.00 } \\
&NAP &51.40  &49.00  &50.30  &\textbf{52.30 } \\
\bottomrule

\end{tabular}}

\caption{Slot-f1 results for Slot Filling on the validation set. Bold entries are the best-performing models for each experiment and setting.}\label{tab: slots}
\end{table}
\noindent{\textbf{Official Shared Task (Test) Results.}} Our findings regarding the performance of larger models are also observed in the test set. Table 6 presents the evaluation results for both slot filling  and intent classification tasks across all three target languages. Our mT0 models strongly outperform the baseline models. Specifically, our mT0 models outperformed the baseline models in all target languages for the intent classification task, highlighting the effectiveness of larger models for intent classification. Moreover, our mT0 models also outperform the baseline models in two of the target languages for slot filling task, further indicating the superiority of larger models for sentence-level classification tasks. The improvement in scores for intent classification is more evident than for slot filling. The larger improvement in scores for intent classification may be correlated with the fact that for our data augmentation experiment on paraphrasing and machine translation, we were only able to augment data for intent classification, resulting in a larger improvement in performance for this task compared to slot filling.

It is worth noting that we use the validation set for model selection, which resulted in higher scores than those achieved on the test set. This is because the validation data is similar to the data used during training, while the test data is entirely new and unseen. As a result, the test scores may be lower due to differences in the distribution of data between the training and test sets. Nevertheless, our mT0 models consistently outperform the baseline models on the test set, providing further evidence for the effectiveness of larger models in SID tasks.

\begin{table}

\centering
 \renewcommand{\arraystretch}{1.1}
\resizebox{\columnwidth}{!}{%
\begin{tabular}{llcc}
\hline
\textbf{Task} & \textbf{Lang.} & \textbf{\begin{tabular}[c]{@{}l@{}}Baseline\end{tabular}} & \textbf{\begin{tabular}[c]{@{}l@{}}mT0-large\end{tabular}} \\ \hline
                     & ST  & $44.61$ & $\textbf{46.41}$ \\
\textbf{Slots}    & GSW & $26.23$ & $\textbf{27.39}$ \\
                     & NAP & $\textbf{48.01}$ & $38.82$ \\ \hline
                     & ST  & $61.00$ & $\textbf{89.40}$ \\
\textbf{Intents} & GSW & $51.67$ & $\textbf{81.60}$ \\
                     & NAP & $61.00$ & $\textbf{85.40}$ \\ \hline
\end{tabular}}
\caption{Results on the test set for both SID tasks. \textbf{Bold} entries indicate the model's performance compared to the baseline model.} \label{tab:test_res}
\end{table}
\section{Conclusion}\label{sec:conc}

We described our contribution to the SID4LR~\cite{2023-findings-vardial} shared tasks. Our models target both the slot and intent sub-task in three proposed low-resource languages,  namely, Swiss German, South Tyrolean, and Neapolitan. We test the utility
of existing pretrained language models such as mT0~\cite{muennighoff2022crosslingual}  on the intent detection and slot filling tasks. We show that such models can lead to improving the results of the baseline with an average of +$27$ F\textsubscript{1} points. In the future, we intend to use mT0 to jointly model the intent detection and slot filling tasks  for improving overall performance.

\section*{Acknowledgements}
We gratefully acknowledge support from the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 435-2018-0576; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Compute Canada (CC),\footnote{\href{https://www.computecanada.ca}{https://www.computecanada.ca}} UBC ARC-Sockeye,\footnote{\href{https://arc.ubc.ca/ubc-arc-sockeye}{https://arc.ubc.ca/ubc-arc-sockeye}} and Advanced Micro Devices, Inc. (AMD). Any opinions, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of NSERC, SSHRC, CFI, CC, AMD, or UBC ARC-Sockeye.

\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}



\end{document}